# Web-CogReasoner: Towards Knowledge-Induced Cognitive Reasoning for Web Agents

### v1

@@ 3.3 Web-CogDataset
3.4 Web-CogBench
To rigorously evaluate the cognitive capabilities fostered by our knowledge-centric framework, we introduce Web-CogBench. While our training dataset is structured by knowledge types (Factual, Conceptual, Procedural), Web-CogBench assesses agent performance through the lens of three corresponding cognitive abilities: Memorizing, Understanding, and Exploring. This benchmark is curated from a representative subset of our Web-CogDataset and is designed to measure how well an agent translates its learned knowledge into practical skills within complex web contexts. Detailed statistics of Web-CogBench are provided in Table 2. The three evaluation dimensions directly mirror our hierarchical knowledge framework:
@@ 4.1 Problem Setup
We model the interaction between the environment and Web-CogReasoner as a partially observable Markov decision process (POMDP): $P=(S,A,O,K,T,R)$, where $S$ denotes the webpage status, $A$ is the action space (See complete operations in Table 1). $O$ is the observation space, and $K$ represents the internal knowledge of Web-CogReasoner. $T:S\times A\rightarrow S$ is the transition function.
@@
$(h_{t},a_{t})=\pi_{\theta}(\cdot|K,I,Q,o_{1},h_{1},a_{1},\dots,o_{t})$
Finally, the overall performance is assessed using a binary reward function $R$, which returns 1 only if the agent successfully completes the user-specified task at the final step, and 0 otherwise.
4.2 Web-CogReasoner Framework
We introduce Web-CogReasoner, a knowledge-driven reasoning system built on LVMs and trained on the Web-CogDataset in Section3.3. It solves complex web tasks by performing chain-of-thought reasoning grounded in knowledge acquired during training. We define web agent actions set based on common user behaviors observed during web browsing. At each interaction step, the model receives a screenshot of the current webpage along with its corresponding AX Tree. The model integrates visual, structural representations with other contextual cues and history trajectories to predict the next action. Once an action is predicted, it is executed by an automated web agent. This process iterates until the task is completed or reach the maximum step limit. The following section details the design of our agent and reasoning framework:
Observation Space
To effectively understand and interact with complex web interfaces, our web agent takes as input both webpage screenshots and corresponding AX Tree. The screenshot provides rich visual context essential for grounding visual reasoning, while the AX Tree offers a structured semantic representation of page elements. Specifically, the AX Tree encodes key properties such as role (indicating the type of an element), name (how the element should be referred to), description (additional descriptive content), and state (e.g., whether a checkbox is checked or a dropdown is expanded).
Action Space
To enable web agents to operate in real-world environments, we define an action space (Table 3) grounded in common user interactions during web browsing. Each action $a_{t}$ operates on a target element identified by a unique ID from the current observation $x_{t}$. For agents using AX Tree representations, this ID refers to a node within the tree.
@@ 4.2 Web-CogReasoner Framework
Knowledge-driven CoT Reasoning
Figure 4: An overview of the Web-CogReasoner framework, illustrating the Knowledge-driven Chain of Thought (CoT) process. The agent initiates with a Task and observes the webpage (Observe Space). The Knowledge-driven COT Reasoning module then systematically leverages different types of Web-CogKnowledge—Factual Knowledge (blue), Conceptual Knowledge (green), and Procedural Knowledge (yellow). This structured reasoning guides the Planning phase, which decomposes the task and formulates a step-by-step strategy. The process culminates in a concrete Action to be executed on the webpage.
To instantiate the hierarchical Web-CogKnowledge Framework in a reasoning-capable agent, we introduce the Web-CogReasoner, a structured cognitive web-agent that generates a Knowledge-driven Chain-of-Thought (CoT) to guide decision-making in web environments. As illustrated in Figure 4, this CoT is not a monolithic or opaque text block; rather, it is a scaffolded reasoning template, where each segment is explicitly grounded in a distinct type of web knowledge. This design ensures interpretability, traceability, and faithful knowledge alignment throughout the reasoning process.
The CoT construction follows a layered progression, with each stage corresponding to a cognitive function defined in the Web-CogKnowledge taxonomy:
  • Factual Knowledge (highlighted in blue) constitutes the foundation of reasoning by anchoring the model in observable page facts. This phase populates the Webpage Layout Description and Key Element Analysis sections. The agent extracts concrete interface features such as DOM structure, ARIA attributes (e.g., role=‘search’), accessible names, and current page states. At this stage, the agent is essentially answering the question: “What is on the page?”
  • Conceptual Knowledge (highlighted in green) builds upon the factual layer by introducing semantic understanding. Within the same sections, the model interprets the purpose of key UI components and predicts their behavior under user interaction. For instance, it explains the functional intent of the search bar (“allows users to input a location”) and anticipates changes to the page layout (“the list of stores will update after input”). This phase reflects the agent’s capacity to comprehend: “What does this element mean?” and “How does it contribute to the task?”
  • Procedural Knowledge (highlighted in yellow) informs goal-oriented planning and task decomposition. It is primarily instantiated in the Task Recap, Task Decomposition, Step-by-Step Reasoning, and Final Action Summary sections. The Reasoner infers the user’s intent from the task prompt, predicts the next sub-goal for interaction (e.g., “Users will type in their desired location”), and generates a sequential plan:
@@
      Enter the ZIP code 90028 into the search bar;

### v2

@@ 4.1 Problem Setup
We model the interaction between Web-CogReasoner and the environment as a partially observable Markov decision process (POMDP): $P=(S,A,O,K,T,R)$, where $S$ denotes the webpage state, $A$ the action space (Table 10), $O$ the observation space, $K$ the internal knowledge, $T$ the transition function, and $R$ the reward function. At each step $t$, the agent receives a screenshot and its accessibility tree (AX Tree), forms a reasoning thought $h_{t}$, and selects an action $a_{t}$ under policy $\pi_{\theta}$. This process continues until task completion or step limits are reached, with binary rewards indicating success or failure.
4.2 Framework Overview
Web-CogReasoner is a knowledge-driven reasoning system built on LVMs and trained on Web-CogDataset (Section 3.3). As illustrated in Figure 3, it tackles complex web tasks by generating a Knowledge-driven Chain-of-Thought (CoT), in which each reasoning stage is explicitly grounded in a layer of Web-CogKnowledge.
Figure 3: Illustrating the Knowledge-driven Chain of Thought (CoT) process.
Knowledge-driven CoT Reasoning
The core of Web-CogReasoner is a structured chain-of-thought (CoT) reasoning process (Figure 3), decomposed into three layers: Factual (identifying page elements and states: “What is on the page?”), Conceptual (inferring roles and interactions: “What does it mean?”), and Procedural (planning goal-directed steps: “How to accomplish the task?”). This layered reasoning maps task prompts to executable actions: $\textbf{Task Prompt}\rightarrow\textbf{Knowledge-driven CoT}\rightarrow\textbf{Plan}\rightarrow\textbf{Action}$. The agent initiates with a Task and observes the webpage (Observation Space). This structured reasoning guides the Planning phase, which decomposes the task and formulates a step-by-step strategy. The process culminates in a concrete Action to be executed on the webpage.
5 Experiment
5.1 Experimental Setup
Models and Baselines
Our proposed Web-CogReasoner builds on Qwen2.5-VL-7B (Bai et al., 2025), extended with a knowledge-induced reasoning framework. Training details are provided in Appendix A.3.1. For comparison, we include diverse baselines: the vanilla Qwen2.5-VL-7B (zero/few-shot), OpenWebVoyager (He et al., 2024b) in IL and Max variants, the open-source UI-TARS-7B-SFT (Qin et al., 2025), and two powerful commercial LVMs, Claude Sonnet 4 and Gemini 2.5 Pro. Together these cover foundational LLMs, end-to-end web agents, and general-purpose multimodal models.
Datasets
Web-CogReasoner is trained via supervised fine-tuning on the curated Web-CogDataset (Sec. 3.3), which aligns samples with Factual, Conceptual, and Procedural knowledge using screenshots, accessibility trees, and reasoning trajectories. Evaluation is conducted on four datasets. Our custom Web-CogBench (Sec. 3.4) assesses cognitive dimensions of Memorizing, Understanding, and Exploring. VisualWebBench (Liu et al., 2024b) provides 1.5K curated tasks across 139 real websites, testing grounding and reasoning in diverse settings. WebVoyager (He et al., 2024b) contains 643 queries across 15 seen sites, measuring performance in familiar environments. Online Multimodal-Mind2Web (Deng et al., 2023) evaluates cross-task and cross-website generalization with queries from both known and unseen domains.
5.2 Main Results
Table 2: Performance evaluation on the Web-CogBench benchmark.
Model Element Attribute Rec Next Page Pre Source Element Pre Element Understanding Claude Sonnet 4 79.7 93.5 62.5 62.8 Gemini 2.5 Pro 79.8 94.6 84.4 62.6 Qwen2.5-VL-7B 53.2 83.9 65.6 60.0 UI-TARs-7B-SFT 63.5 88.0 31.3 48.0 Web-CogReasoner (Ours) 91.4 93.5 87.5 69.2
WebPage Understanding User Intent Pre Popup Close Single Step Exp Overall
Claude Sonnet 4 54.3 64.7 100 96.8 76.8
Gemini 2.5 Pro 73.5 51.9 96.6 98.4 80.4
Qwen2.5-VL-7B 62.0 51.9 91.4 90.3 69.8
UI-TARs-7B-SFT 48.0 32.4 25.9 33.9 46.4
Web-CogReasoner (Ours) 79.0 61.4 98.3 95.2 84.4
Results on Web-CogBench
We evaluate Web-CogReasoner on Web-CogBench, which assesses foundational knowledge and cognitive reasoning across twelve web tasks. As Table 2 shows, our model outperforms both commercial and open-source baselines, owing to the integration of structured Web-CogKnowledge (factual, conceptual, procedural) with Knowledge-driven Chain-of-Thought reasoning. This combination enables accurate perception of web elements and informed, step-wise decision-making. The strong synergy between visual recognition and cognitive reasoning allows Web-CogReasoner to excel on diverse tasks and generalize effectively to web navigation scenarios. We also provide a detailed LVM evaluation protocol and inter-rater reliability analysis in Appendix A.4 to validate the robustness of our scoring metrics.
Table 3: Performance evaluation on the VisualWebBench benchmark.
Model Perception-Oriented Tasks Perception Avg Reasoning-Oriented Tasks Reasoning Avg Overall Avg
WebQA HeadOCR OCR Element Ground Action Prediction Action Ground
Claude Sonnet 4 73.3 72.6 96.2 80.7 81.1 96.1 96.3 91.2 85.9
Gemini 2.5 Pro 74.9 70.8 95.1 80.3 91.8 96.8 90.3 93.0 86.6
Qwen2.5-VL-7B 70.8 71.7 81.4 74.6 77.5 86.8 68.0 77.4 76.0
UI-TARs-7B-SFT 71.3 78.7 97.2 82.4 91.8 91.8 85.4 89.7 86.0
Web-CogReasoner (Ours) 67.2 72.6 97.0 79.0 96.4 96.1 88.4 93.6 86.3
Results on VisualWebBench
We evaluate visual understanding on VisualWebBench (Liu et al., 2024b). As Table 3 shows, Web-CogReasoner achieves the highest average score (86.3%), slightly above UI-TARs (86.0%). However, UI-TARs performs poorly on Web-CogBench (46.4%), highlighting that strong visual perception alone does not ensure robust cognitive reasoning. In contrast, Web-CogReasoner consistently excels across both visual and reasoning benchmarks, demonstrating the effective integration of precise visual perception with structured knowledge-driven reasoning — a dual capability essential for reliable web agents.
Table 4: Task success rates on the WebVoyager. The ”Overall” score is the average success rate.
Agent Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
Claude Sonnet 4 26.7% 87.8% 48.8% 69.8% 68.3% 2.3% 45.5% 83.3%
Gemini 2.5 Pro 60.0% 63.4% 62.8% 67.4% 68.3% 9.1% 56.8% 73.8%
Qwen2.5-VL-7B 0.0% 0.0% 0.0% 4.7% 0.0% 0.0% 0.0% 2.3%
@@ 5.2 Main Results
[…]
BBC News Cambridge Dictionary Google Flights Google Map Huggingface Wolfram Alpha Overall
Claude Sonnet 4 23.8% 37.2% 4.8% 80.5% 48.8% 82.6% 47.7%
Gemini 2.5 Pro 52.3% 76.7% 4.8% 75.6% 58.1% 82.6% 54.9%
Qwen2.5-VL-7B 0.0% 11.6% 0.0% 2.4% 7.0% 2.2% 2.2%
[…]
Web-CogReasoner (Ours) 14.3% 55.8% 9.5% 39.0% 37.2% 39.1% 30.2%
Table 5: Performance comparison on the Online-Mind2Web under cross-task and cross-websites
Agent Cross-task (Unseen Task) Cross-web (Unseen Websites)
Claude Sonnet 4 44.9% 35.3% 40.0% 40.2% 45.5% 6.7% 14.0% 21.7%
Gemini 2.5 Pro 46.9% 35.3% 28.3% 37.5% 42.4% 10.0% 23.3% 25.5%
OpenWebVoyagerMax 22.4% 29.4% 15.2% 20.5% 3.0% 8.7% 23.3% 11.7%
Qwen2.5-VL-7B 2.2% 0.0% 0.0% 1.0% 3.0% 0.0% 0.0% 1.0%
OpenWebVoyagerIL 8.2% 5.9% 4.3% 6.3% 3.0% 5.8% 4.7% 6.6%
Web-CogReasoner (Ours) 16.3% 23.5% 15.2% 17.0% 12.1% 7.7% 9.3% 10.1%
Results on Online Web Tasks
[…]
Table 6: Average steps per successful task across different benchmarks.
Agent WebVoyager Mind2Web Cross-Task Mind2Web Cross-Web Final Avg
Claude 7.35 10.89 11.04 9.76
Gemini 6.68 7.74 10.30 8.24
[…]
Qwen2.5-VL-7B 7.69 12.00 13.00 10.9
[…]
Ours 4.73 7.37 8.89 7.00
Results on Average Steps
Table 6 reports the average number of steps for successful online tasks. Our approach consistently achieves high efficiency, particularly in cross-domain scenarios, indicating that the model effectively balances streamlined task execution with robust generalization to unseen environments.
5.3 Ablation Study
To empirically validate the effectiveness of our Web-CogKnowledge Framework, we conduct a two-fold ablation study. First, we evaluate the cumulative gains of our curriculum learning strategy (Table 7). Second, to address specific inquiries regarding the necessity of each knowledge layer and the reasoning mechanism, we provide a detailed component analysis on both Web-CogBench and WebVoyager (Tables 8 and 9).
Table 7: Cumulative Gains: Impact of progressive knowledge training on Web-CogBench.
Model Configuration Memorizing Understanding Exploring Overall
Qwen2.5-VL-7B (Base Model) 67.6 61.0 77.9 69.8
+ Factual Knowledge (S1) 85.5 (+17.9) 64.2 60.1 72.1
+ Conceptual Knowledge (S2) 88.1 75.5 (+11.3) 65.8 78.3
+ Procedural Knowledge (S3) 90.8 74.1 85.0 (+19.2) 84.4
Cumulative Impact of Curriculum Learning
We evaluate the roles of Factual, Conceptual, and Procedural knowledge through an ablation study that incrementally augments the base Qwen2.5-VL-7B model and measures performance on Web-CogBench. Factual Knowledge provides the perceptual grounding needed for accurate element recognition, strengthening the Memorizing dimension. Conceptual Knowledge introduces semantic structure and functional understanding, improving the Understanding dimension and enabling basic multi-step behaviors. Procedural Knowledge adds goal-directed planning and execution, yielding major gains in the Exploring dimension. Qualitative examples illustrating how perception, interpretation, and planning evolve across stages are presented in Appendix A.5. Together, the results show that each knowledge layer is essential, and their integration is key to Web-CogReasoner’s cognitive robustness and performance.
Table 8: Component Analysis on Web-CogBench: Validating hierarchical dependency.
Model Memorizing Understanding Exploring Overall
Base model 67.6 61.0 77.9 69.81
S1 only 85.5 64.2 60.1 72.12
S2 only 59.88 68.03 60.00 61.96
S3 only 52.82 46.40 78.00 60.66
S1+S2 88.1 75.5 65.8 78.33
S1+S3 85.11 53.53 82.31 76.17
S2+S3 64.87 69.74 81.41 72.29
S1+S2+S3 (Full) 90.8 74.1 85.0 84.44
Table 9: Impact of Knowledge & KCoT on WebVoyager, real-world online tasks.
Model Amazon Cambridge Coursera GitHub Overall
S1 only 12.19% 25.58% 14.28% 7.14% 12.67%
S3 only 17.07% 11.62% 16.66% 14.28% 13.14%
S1+S3 29.26% 34.88% 28.57% 16.66% 23.47%
S1+S2+S3 (w/o KCoT) 19.51% 51.16% 26.19% 23.80% 25.35%
S1+S2+S3 (w/ KCoT) 31.7% 55.8% 54.8% 29.3% 42.9%
Hierarchical Dependency of Knowledge
To demonstrate that our stages are not merely isolated skills but hierarchically dependent layers, we present a detailed breakdown in Table 8 and Table 9.
  • Low-level Knowledge is a Prerequisite: As seen in Table 8, single-stage models (S2 only, S3 only) perform poorly on comprehensive metrics. Crucially, adding Factual training (S1) significantly boosts the performance of higher-level stages. For example, on WebVoyager (Table 9), combining S1 with S3 nearly doubles the success rate compared to S3 alone (23.47% vs. 13.14%). This proves that procedural exploration (S3) cannot function effectively without accurate factual grounding (S1).
  • Integration is Critical: While specific stages excel at their corresponding metrics (e.g., S3 on Exploring), only the fully integrated model (S1+S2+S3) achieves robust performance across all dimensions, confirming that complex web agents require a complete cognitive stack.
Reasoning Activation via KCoT
Finally, we investigate the role of our reasoning framework. While the full combination of data (S1+S2+S3) builds a strong latent representation, explicit reasoning is required to utilize it. As shown in Table 9, removing the Knowledge-driven Chain-of-Thought (w/o KCoT) causes a sharp drop in online success rate from 42.9% to 25.35%. This indicates that KCoT acts as a crucial activator, bridging the gap between possessing knowledge and applying it dynamically for decision-making.
6 Conclusion
We present Web-CogReasoner, a cognitive-inspired framework for web agents that systematically instills Factual, Conceptual, and Procedural Knowledge, following Bloom’s Taxonomy. By coupling the Web-CogKnowledge Framework with the Web-CogDataset and Web-CogBench, our approach enables interpretable, step-wise reasoning through knowledge-driven Chain-of-Thoughts, yielding strong performance on complex web navigation and instruction-following tasks. Ablation and qualitative analyses confirm the indispensability of each knowledge stage, demonstrating how a structured curriculum produces robust perceptual and cognitive capabilities. While current results rely on imitation learning, future work aims to integrate reinforcement learning to enhance exploration, generalization, and autonomous discovery of procedural knowledge, advancing toward truly adaptive and self-directed web agents.
Appendix A Appendix
A.1 Web Cognition and Data Details
A.1.1 Bloom’s Taxonomy
  • 1.
  • 2.
  • 3.
    Procedural Knowledge: The knowledge of how to perform a task or inquiry. This involves an understanding of specific skills, algorithms, techniques, and methods, representing a shift from ”knowing-what” to ”knowing-how.”
@@ A.1.1 Bloom’s Taxonomy
A.1.2 Action Space Overview
Table 10 summarizes the action space of Web-CogReasoner for web interaction tasks. The table categorizes actions into two groups: (1) element-specific operations, such as clicking, typing, double-clicking, and scrolling individual elements; and (2) page-level control actions, including navigation commands, task restart, waiting, and final answer submission. This structured action space enables the agent to perform a diverse set of interactions, facilitating comprehensive exploration and manipulation of web pages in a controlled and systematic manner.
A.1.3 Stages of Human Knowledge and Cognitive Development
A.1.4 Web data introduction
Figure 5 presents the statistics of the selected websites, grouped by category. The visualization highlights the relative proportions of different types of websites included in our study, providing an overview of the dataset composition. Such a distribution analysis is crucial for understanding potential biases, ensuring coverage across diverse web domains, and evaluating the generalizability of models trained or tested on these websites.
Figure 5: Statistics of selected websites by category.
Data Composition and Balance
To avoid domain overfitting and ensure both interaction depth and broad generalization, Web-CogDataset employs a strategic hybrid data composition:
@@ A.1.4 Web data introduction
    Depth via Self-Collected Data: We selected 14 complex websites for high-depth interaction mining to capture intricate logic often missed by general crawls. As shown in Figure 5, we strictly maintained category balance across E-commerce, Finance, Developer Tools, and Social Media within this subset to prevent bias toward any single domain.